If I want a patio that has 9 times the are by what factor do I have to increase the side?
I think that the answer is 3. I know that if I increase x by a factor of t, then f is increasedc by a factor of t^k. It is doing is backwards that I have a problem with. So my idea is that if f is 9 times its original size, then because we are dealing with area the k is 2. which means that 9=x^2. which would make my answer 3. Right?

Answer by edjones(8007)

L*w=A
9L*w=9(L*w)=9A
You have to increase it by a factor of 9.
Ed